Science Report – Nov 6th

Crew 214 Science Report 6 November 2019
SOL: 10
Crew 214 Ð Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The seventh scientific EVA for the mission saw us return to Tank Wash #2 to collect 23 samples from the micrometeorite grid. Four of these had been compromised and it remains to be seen how this will affect the study. 64% of the plots have now been sampled.

With the disappointing failure of the regolith heater prior to arrival at MDRS, a modified plan to demonstrate water extraction was devised. Instead of an in situ test of the device, the gypsum and gypsiferous soil samples collected yesterday would undergo the process in the Science Lab. After documenting the samples, 6 small subsamples of ground gypsum were heated. Within minutes of each subsample being heated, water vapour was observed and video recorded. Further testing of the gypsiferous soil will be attempted tomorrow.

Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from yesterday’s EVA were processed and experimental procedures commenced.

Closed Loop Food Waste Study: The day’s waste for Sol 10 is: 51 grams; consisting of 175 calories, 4g fat, 2g saturated fat, 328mg sodium, 26g carbohydrates, 3g fibre, 8g sugar, 10g protein.

EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.

EVA Suit Maintenance: Nominal

Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

Science Report – November 05th

Crew 214 Science Report 5 November 2019
SOL: 9
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The sixth scientific EVA for the mission saw us travel to White Moon (517750E 4254520N) at the junction of Brahe Highway and Cow Dung Road in search of hydrated minerals. Immediately on arrival, gypsum crystals (CaSO4.2H2O) were observed protruding from the regolith and we were able to collect two soil samples and 9 rock samples. From these we intend to demonstrate the liberation of water through heating from both the crystals and the soil.
On our return to the the hab, we called into the Tank Wash #2 site and collected an additional 3 samples for the micrometeorite search. This is 43% of the planned samples. Unfortunately, someone has visited the grid, trampled through a number of plots and compromised the integrity of the experiment.
Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from the morning’s EVA are awaiting processing before undergoing experimental procedures within the science dome.
Closed Loop Food Waste Study: The day’s waste for Sol 9 is: 25 grams; consisting of 127 calories, 1g fat, <1g saturated fat, 69mg sodium, 70g carbohydrates, <1g fibre, 2g sugar, 5g protein.
EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.
EVA Suit Maintenance: Nominal
Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

Science Report – November 4th

Crew 214 Science Report 4 November 2019
SOL: 8
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The fifth scientific EVA for the mission saw us returning to Tank Wash #2 and collecting 28 additional samples for the micrometeorite search. This is 40% of the planned samples.

Green Hab: Following the commencement of planting within the green hab, access through the tunnel between the main hab and the green hab has been opened. The plants were watered with approximately 3.5 litres of water, with a remaining 4.5 litres remaining in the watering can. The plant soil appeared fairly moist, so it is suggested crew refrain from over watering the new plants.

Science Dome: Samples returned from the morning’s EVA have now been processed within the science dome and limited observations of material recovered through the binocular microscope have been carried out.

Closed Loop Food Waste Study: The day’s waste for Sol 8 is: 51 grams; consisting of 159 calories, 3g fat, >1g saturated fat, 199mg sodium, 42g carbohydrates, 3g fibre, 13g sugar, 6g protein. Extra food supplies were received and added to the baseline stock database.

EVA: Preparations for an EVA to White Moon at the junction between Brahe Highway and Cow Dung Road have been carried out.

EVA Suit Maintenance: Images of helmet 5 have been forwarded to CapCom detailing the partial failure, the battery to Suit 10 has been examined and it is the suit itself that is faulty.

Additional Activities: Nil.

Science Report Nov 3rd

Crew 214 Science Report 3 November 2019

SOL: 7

Crew 214 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Operations: No activity during programmed day off other than the closed loop food waste study, where daily tallies continued and end-of-rotation preparations commenced.

Closed Loop Food Waste Study: The day’s waste for Sol 7 is: 45 grams; consisting of 164 calories, 5g fat, 4g saturated fat, 159mg sodium, 21g carbohydrates, 2g fibre, 1g sugar, 24g protein.

Additional Activities: Nil.

Science Report – November 2nd

Crew 214 Science Report 2Nov2019
SOL: 6
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: Preliminary viewing, through the binocular microscope, of the particles recovered from yesterday’s EVA clearly showed the morphology of the particles and that some were spherical, however positive identification of micrometeorites could not be confirmed.
The fourth of the scientific EVAs was scheduled for the afternoon work period. This involved driving to site #2 at Tank Wash and continuing sample collection from the 1m x 1m plots. During the 2 hour EVA, 10 samples were collected and all appear to contain magnetic material.
On the return to the hab, the Cactus Road #1 site was revisited to allow another crew member an opportunity to operate the heliograph. With line of sight radio communications established, the heliograph was successfully operated and video records collected.
Tabulation of the newly arrived food stock macro and micronutrients for the waste study was completed.
Closed Loop Food Waste Study: The day’s waste for Sol 6 is: 15 grams; consisting of 44 calories, 1g fat, >1g saturated fat, 16mg sodium, 7g carbohydrates, 1g fibre, >1g sugar, 6g protein. At the end of the first week of the rotation, over 613 grams (2791 calories) of food have been wasted at a rate of approximately 400 calories per day, or 10 grams per person per day. This waste was proportionally derived from carbohydrates (63%), protein (15%) and fats (22%), which was a slight shift from the macronutrient ratios of crew stock inventory from the start of the rotation (C58%, P12%, F31%).
Additional Activities: Tidying and cleaning chores around the hab.

Science Report – November 1st

Crew 214 Science Report 1 November 2019
SOL: 5
Crew 214 Ð Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The third of the scientific EVAs was scheduled for the afternoon work period. This involved driving to site #2 at Tank Wash and continuing setting out the 100 square metre (100 1m x 1m plots). During the 2 hour EVA, setting out of the grid was completed and the first samples collected. This involved hovering the heavy duty magnet just above the ground surface and systematically ‘sweeping’ it across the 1 square metre of the plot. Two plots were sampled and very fine magnetic particles were recovered from both plots!
On the return to the hab, the Cactus Road #1 site was revisited to allow another crew member an opportunity to operate the heliograph. With line of sight radio communications established, the heliograph was successfully operated and video records collected.
Tabulation of the baseline food stock macro and micronutrients for the waste study is being updated with the arrival of further supplies. Closed Loop Food Waste Study: The day’s waste for Sol 5 is: 108 grams; consisting of 383 calories, 6g fat, 3g saturated fat, 176mg sodium, 53g carbohydrates, 4g fibre, 3g sugar, 24g protein. The bread production errors were turned into tasty crackers/cookies (#nowaste).
Additional Activities: Nil.

Science Report – Oct 31st

Crew 214 Science Report 31 October 2019
SOL: 4
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: The second of the scientific EVAs was scheduled for the morning work period. This involved driving to site #2 at Tank Wash and establishing a 100 square metre (100 1m x 1m plots) on a relatively undisturbed natural surface. During the 2 hour EVA, the outline of the grid was established and three east-west delineators positioned. The task is incomplete.

A secondary experiment involved testing the magnetism of material brought to the surface at an adjacent anthill. Though it was hypothesized that this material might contaminate the sample plots; there was zero evidence of magnetic material at the surface.

A tertiary experiment was conducted at the Cactus Road #1 site before the return to the hab. With line of sight communication possible, a heliograph was used to demonstrate an emergency communications method should voice contact fail. The demonstration was successful and video records collected.

Tabulation of the baseline food stock macro and micronutrients for the waste study was completed.

Closed Loop Food Waste Study: The day’s waste for Sol 4 is: 229 grams; consisting of 891calories, 15g fat, 9g saturated fat, 1533mg sodium, 126g carbohydrates, 8g fibre, 4g sugar, 42g protein. Waste was significantly larger due to bread production errors. Otherwise, less waste is generally produced at breakfast, with almost all occurring at lunch or dinner, depending on the type of food cooked. The majority is preparation waste with no spoilage waste to date and a small amount of plate waste.

Additional Activities: Reinforcing the coverings for the tunnels continued throughout the day as time became available.

Science Report – October 30th

Crew 214 Science Report 30 October 2019
SOL: 3
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The first of the scientific EVAs was scheduled for the afternoon work period. This involved a 2 hour traverse in the rovers to Tank Wash and south of the Galileo and Cactus Road junctions. At both these locations, relatively flat relatively undisturbed areas were identified and GPS surveyed and photographed so that a 10 square metre grid could be deployed from which samples targeting mircometeorites would be collected.
In preparation for the sampling, during the morning work period, the Science Dome’s large open area was utilized to construct the 1m x 1m grid square boundaries in readiness for deployment.
Concurrent with the preparations for micrometeorite samplings and EVA, tabulation of the baseline food stock macro and micronutrients was undertaken for the waste study.
Closed Loop Food Waste Study: The day’s waste for Sol 3 is: 89 grams; consisting of 293 calories, 1g fat, >1g saturated fat, 555mg sodium, 65g carbohydrates, 6g fibre, 31g sugar, 10g protein.
Additional Activities: Reinforcing of the coverings for the tunnels was undertaken throughout the day as time became available.

Science Report – October 29th

Crew 214 Science Report 29 October 2019
SOL: 2
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler

Science Operations: Weather conditions prevented EV activities during the morning, so all crew underwent instruction in donning and doffing both the exploration and two piece suits. In the afternoon, EVAs were approved and two short training activities were carried out. All crew have now become familiar with the operation of both types of suits, hab egress and ingress, operations with rovers and mobility on the surface.

Closed Loop Food Waste Study: The day’s waste for Sol 2 is: 89 grams; consisting of 339 calories, 10g fat, 5g saturated fat, 1619mg sodium, 56g carbohydrates, 3g fibre, 10g sugar, 5g protein.

Science Report Oct 27th

Crew 214 Science Report 27-Oct-2019

Submitted by Science Officer Andrew Wheeler

1. EVA Suit Maintenance: The hab has been set up following arrival. All backpacks have been plugged in and fully charged. Helmets are clean and contained in their cloth bags except for one helmet, for which, the cloth bag is missing. Radios have been tested and are fully charged. Communication earpieces remain to be tested for faulty performance.

2. Green Hab: The green hab has been activated and is awaiting water and additional soil base before sowing of this seasons crops.

3. Astronomy Dome Operations: No astronomical observations have been planned for this crew rotation.

4. Science Dome Operations: The Science Dome has been activated and is awaiting EVAs for sample acquisition and sample testing.

5. RAM Operations: The RAM Dome has been activated and is ready for contingencies.

6. EVA: No preparations or activities have been undertaken to date. The access routes have been identified to allow sample locations to be identified.

7. Closed Loop Food Waste Study: This study has been initiated and is ongoing. Day’s waste: 79g; consisting of 323 calories, 10g fat, 5g saturated fat, 1179mg sodium, 52g carbohydrates, 7g fibre, 6g sugar, 7g protein.

8. Additional activities: Nil